Amp nucleosidase gives a rate enhancement of 6.0 x 1012 over the nonenzymatic reaction. What is the ∆∆G of this enzyme?

A.) 9.0 kJ/mol

B.)75.8 kJ/mol

C.) 60 kJ/mol

D.) 49 kJ/mol
Answer #1

The complete step wise calculation of the change in Gibb's free energy is given as below:-

It is rate given that, Constant (k) = 6.0x1012 Gas Constant (R) = 8.3148 J/molik T= 25 °C=(25+273) * = 298k IKJ=1000 J. DG °

AGT = +RT lnk - =(+8.314 Jlmolok) x (2986) In (6.0x10/2) = + 2477.572 (log 6 +12) x2-303 = 5705.85 X 12.778 J/mol I = 72.909